Edgar is 36 years older then Bob
Simone is 20 years younger then Bob
The today of the 3 ages if 136
How old is bob?

Found 3 solutions by Boreal, josgarithmetic, Alan3354:
Answer by Boreal(15235)

Bob is x
Edgar is x+36
Simone is x-20
They add to 136
so 3x+16=136
3x=120
x=40 Bob
Edgar is 76
Simone is 20

Answer by josgarithmetic(39620)

--------------
--------------
Edgar is 36 years older then Bob
Simone is 20 years younger then Bob
The today of the 3 ages if 136
How old is bob?
--------------
--------------

The what of the three ages? "TODAY"?
PERSON          AGE
EDGAR           b+36
BOB              b
SIMONE          b-20
The what?        136

You probably meant, "total age now"

Answer by Alan3354(69443)

Edgar is 36 years older then Bob
Simone is 20 years younger then Bob
The today --total?-- of the 3 ages if 136
How old is bob?
----------
E = B + 36
S = B - 20
===================
E + S + B = 136
etc